# Comments on “Exploring Quantum Statistics for Dirac and Majorana Neutrinos using Spinor-Helicity techniques” (arXiv:2507.07180 [hep-ph])

### v1

@@ Abstract
We give our comments on Ref. [1] which critiques our idea of exploring quantum statistics to distinguish between Dirac and Majorana neutrinos proposed in some of our earlier works. We show that the ad-hoc symmetrization of the Dirac case amplitude square advocated in Eqs. (16) and (35) of [1], have no physical basis and it leads to violation of lepton number in the the standard model. Therefore, we consider this symmetrization ‘by hand’ as incorrect and misleading. Moreover, we show that it is even unnecessary for the proof of pDMCT in context of final states containing a neutrino and antineutrino pair.
I What is the primary issue and where does it appear in [1]?
In Ref. [1] the authors aim to provide a critique of our papers [2, 3, 4] where we have explored the possibility whether quantum statistics could help in distinguishing Dirac and Majorana neutrinos. We note that in Ref. [1] the authors have tried to force agreement with the practical Dirac Majorana confusion theorem (pDMCT), while in our works [2, 3, 4] we have explored the domain of applicability of pDMCT in addition to providing a general proof of it in context of final states having neutrino antineutrino pair. We have used the standard quantum field theory in our papers [2, 3, 4] and our approach very well accommodates the previously existing literature.
The underlying basis of the main claim of Ref. [1] where they disagree with our papers, is contained within their Eqs. (33)-(38) of Sec. IV. While discussing the case of Dirac neutrinos, the authors of [1] contend that for the unobserved final states of neutrino and antineutrino one should symmetrize the amplitude squared with respect to the 4-momenta of the neutrino and antineutrino. Just below Eq. (34) of [1] they state that “To make a fair comparison with the Majorana case, as the neutrinos are not observed, it is imperative that both momentum assignments be the anti-neutrino are summed to construct the Dirac matrix-element squared”. This notion of symmetrizing at the probability level for Dirac neutrinos affects the inferences made in Ref. [1]. As explained below, we find such symmetrization “by hand” to be incorrect and misleading. Our comments below exclusively discuss all aspects of this particular issue found in Ref. [1]. It should be mentioned that although we primarily focus on their Sec. IV where they have critiqued our work on the decay [2], the above mentioned issue can also be found in Sec. III of [1] around Eqs. (14)-(16). Finally, we also point out the misinterpretation of the helicity amplitudes given in Eq. (28) of [3] by the authors of [1] in their Eqs. (39)-(43). We clarify these aspects as well.
II Our comments
Our comments are as follows.
@@ II Our comments
Detection and identification of a particle in a detector does not affect the amplitude square which is a purely theoretical computation where one assumes that all the 4-momenta of all the particles are well known. The quantity which gets affected by experimental detection or observation or measurement is the “observable” (such as some asymmetry, or distribution, or differential decay rate, or scattering cross-section etc.) which involves both the amplitude square as well as the phase space considerations. Non-observation of any final particle is well taken care of by suitably performing corresponding phase space integration. We will later show how such phase space integration for invisible neutrinos, without any ‘by hand’ symmetrization of the Dirac amplitude square, directly leads to pDMCT which we have already provided in [2]. But before that we point out how such symmetrization is incorrect and misleading.
@@
In the standard model (SM) the Dirac neutrino () and antineutrino () carry lepton numbers as they are produced via different weak charged currents (CCs) and for , to ensure lepton number conservation within the SM. When considering Dirac neutrino antineutrino pair production via two CCs, as it happens in the decays considered in Ref. [1] as well as in our work [2], the neutrino and antineutrino 4-momenta are paired with the corresponding charged lepton so that lepton number is conserved. This is not required for Majorana neutrinos, in which and the lepton number is not a good (conserved) quantum number. However, adding a term to the Dirac case amplitude square where the neutrino and antineutrino 4-momenta have been exchanged, as done in the right sides of Eqs. (16) and (35) of Ref. [1], amounts to considering and both of which violate lepton number. When working with Dirac neutrinos in the framework of the SM alone without consideration of any new physics contributions, as is the case in decay considered in both [1] and [2], one should avoid considering lepton number violation effects.
3.
@@
III Explanations regarding non-symmetrization of Dirac amplitude square and pDMCT
Here we explain how proving pDMCT does not require the ad-hoc symmetrization of Dirac case amplitude square as done in [1]. If there is no way to infer the energies or momenta of the neutrino and antineutrino experimentally, then any experimental observable, say for Dirac (D) or Majorana (M) neutrinos, would always involve integration over the full 4-momenta of neutrino and antineutrino. In such a case it does not matter whether we assign and to antineutrino and neutrino respectively or vice versa, because
Here act as dummy variables for integration as they both have same integration ranges. In the Majorana case, for the observable we get,
which implies that the observational difference between Dirac and Majorana case observables in such a case is proportional to ,
because the interference term here is always (except in the case of collinear neutrinos, i.e. for ) proportional to as it involves helicity flips. This is the most general way to prove pDMCT in context of processes that have a neutrino and antineutrino pair in their final states. Most importantly,
- 1.
2.
it does not require any by hand symmetrization of the Dirac case amplitude square as done in [1].
One of the most direct violations of pDMCT occurs when one considers the collinear neutrino antineutrino pair, (say). In such a case, considering the decay , it can be shown that Majorana amplitude vanishes in this case due to anti-symmetrization, the amplitude for the Dirac case is non-zero, i.e.
and as given in [2] for Dirac case we get
In Dirac case the probability is not only non-zero, but is also not proportional to , which defies the logic used in Ref. [1]. In fact it remains the same and is non-zero even if one wrongly symmetrizes the amplitude squared as done in [1].
Since quantum statistics of identical particles does not depend on the mass of the particle but its spin, one expects that there would be certain signatures of quantum statistics that explicitly shows this behavior, i.e. there would be some predictions from quantum statistical consideration which do not depend on mass. Our exploration of quantum statistics to distinguish between Dirac and Majorana neutrinos is based on this positive expectation and it is bolstered by the collinear case mentioned above although experimentally we do not yet have a way to study the collinear case. Moreover, it would be unwise to interpolate results of any particular process to a different process which might involve different interactions altogether. For example, in our study of (see Eq. (1) of [2], Eq. (16) of [3], Eq. (21) of [5]) we find that in the SM we always get
which is in confirmation of pDMCT even at the amplitude square level. In comparison, for doubly weak charged current mediated processes such as , when one does not use any ad-hoc symmetrization of Dirac amplitude square, we get from Eqs. (31), (32) of [2],
Then the challenge here is to find out an experimentally measurable quantity, a suitable observable, without completely destroying the finite difference between Dirac and Majorana cases. Our study has helped us realize that there are three major ways one might circumvent pDMCT,
1.
if direct and exchange amplitudes in Majorana case have non-trivially different contributions (Eq. (6) of [2] which gets satisfied in Eqs.(31) and (32)),
- 2.
- 3.
IV Conclusion
We have considered the claims made in Ref. [1] in context of our papers [2, 3, 4] and presented clarifications for considering them to be both invalid and incorrect. The main crux of the arguments presented in our papers [2, 3, 4] is to show that quantum statistics can not only distinguish between Majorana or Dirac types, but also define the domain of validity of pDMCT in the most general way possible in the absence of a formal proof of it until now. As everyone agrees, pDMCT is valid when we integrate over the momenta of neutrino and antineutrino (using the Dirac nomenclature). The proposed ad-hoc symmetrisation in [1] at the probability level for Dirac case simply achieves this by a different procedure, but as explained before we do not consider this approach to be correct. Nevertheless, our main focus in [2, 3, 4, 5, 6] is on the possible exceptions to pDMCT when we can fix the neutrino-antineutrino momenta either directly (possible in principle but difficult experimentally) or indirectly (as pointed out in [2] under special kinematic conditions) or when new physics effects are taken into consideration.
@@ Acknowledgements.
The work of CSK is supported by grants from the National Research Foundation (NRF) of the Korean government (RS-2022-NR070836).References
- [1] I. Bigaran, S. J. Parke and P. Pasquini, “Exploring Quantum Statistics for Dirac and Majorana Neutrinos using Spinor-Helicity techniques,” [arXiv:2507.07180 [hep-ph]].

### v2

@@ Abstract
We give our comments on Ref. [1] which critiques the idea of using quantum statistics to distinguish between Dirac and Majorana neutrinos proposed in some of our earlier works [2, 3, 4]. The ad hoc symmetrization of the Dirac case amplitude square advocated in Eqs. (16) and (35) of [1] has no physical basis and it leads to violation of lepton number in the standard model for Dirac neutrinos. Therefore, this symmetrization ‘by hand’ is in principle incorrect.
I Note Added after publication of [1]
In the published version of the paper [1] we find essentially the same mistakes which we had pointed out before (see Sec. II for a concise version of our original comments).
@@ I Note Added after publication of [1]
Their Dirac case amplitude square for the doubly weak charged decay of the meson: , as given in their Eq. (35) has two terms. The second term has been added by hand to their Eq. (34). The first and second terms of their Eq. (35) are related to each other via exchange of the neutrino antineutrino 4-momenta, . This makes their Dirac case amplitude square symmetric under , even though Dirac neutrino and antineutrino are distinguishable particles. The addition of the exchanged second term without simultaneously interchanging the muon 4-momenta, ie. , implies that the lepton number is violated for the second term of their Eq. (35). The anti-symmetrization of amplitude for Majorana neutrino is enforced by Pauli principle due to the identical nature of Majorana neutrino and antineutrino, and it is unrelated to detection or non-detection of the final neutrinos. Therefore, the ad hoc symmetrization of Dirac case amplitude square under exchange in Eq. (35) of [1] is incorrect.
2.
The Eqs. (16) and (35) of [1] seem to suggest that if one does not observe some of the non-identical final state particles, then one needs to symmetrize the amplitude square of the process by adding terms that exchange the unobserved 4-momenta. We find this as a rather strange claim. It is a well known fact that in any collider experiment some particles escape detection or identification, when they are either very close to the beam pipe or when the tracks get too crowded or if a part of the detector malfunctions or if many long-lived particles are present in the final state which decay outside the detector, etc. In such cases the amplitude square of the process is never symmetrized for non-detection of any of the distinct particles. Instead, the amplitude and amplitude square of the process are calculated assuming that all the non-identical particles are detectable and distinguishable in principle, and the non-detection of any particle is taken care of by appropriately constraining the phase space integral.
3.
In our paper [2], we have used a simplified notation to denote the amplitude of the doubly weak charged decay of the meson: . Our simplified notation for the Dirac case amplitude, as given in Eq. (2) of our paper [2] has the following expanded form,
where we have now explicitly mentioned all the 4-momenta and spins of the final particles. For brevity of expressions we chose to highlight only the dependence, while other dependencies were implicitly always there. As mentioned below our Eq. (3) in [2], the amplitude squares, e.g. , include summation over all final spins and averaging of initial spins. Note that since the parent particle has spin , the average over initial spins is trivial/meaningless in our case.
4.
Just above Eq. (43) of [1], the authors claimed that “… exactly vanishes when … .” This sentence is absolutely misleading. As explained above, the spin information is implicit in our simplified amplitude notation. Therefore, the correct statement would have been “… exactly vanishes when and … .”
II Summary of our Original Comments
In Ref. [1] the authors aim to provide a critique of our papers [2, 3, 4] where we have explored the possibility whether quantum statistics could help distinguish Dirac and Majorana neutrinos. We note that in Ref. [1] the authors have tried to force agreement with the practical Dirac Majorana confusion theorem (pDMCT), while in our works [2, 3, 4] we have explored the domain of applicability of pDMCT in addition to providing a general proof of it in context of final states having neutrino antineutrino pair. We have used the standard quantum field theory in our papers [2, 3, 4] and our approach very well accommodates the previously existing literature. Our comments are as follows.
1.
The underlying basis of the main claim of Ref. [1] where they disagree with our papers, is contained within their Eqs. (33)-(38) of Sec. IV, as well as in Eqs. (14)-(16) of their Sec. III. While discussing the case of Dirac neutrinos, the authors of [1] contend that for the unobserved final states of neutrino and antineutrino one should symmetrize the amplitude squared with respect to the 4-momenta of the neutrino and antineutrino. Just below Eq. (34) of [1] they state that “To make a fair comparison with the Majorana case, as the neutrinos are not observed, it is imperative that both momentum assignments be the anti-neutrino are summed to construct the Dirac matrix-element squared”. This notion of symmetrizing at the probability level for Dirac neutrinos affects the inferences made in Ref. [1].
@@ II Summary of our Original Comments
Detection and identification of a particle in a detector does not affect the amplitude square which is a purely theoretical computation where one assumes that all the 4-momenta of all the particles are well known. The quantity which gets affected by experimental detection or observation or measurement is the “observable” (such as some asymmetry, or distribution, or differential decay rate, or scattering cross-section etc.) which involves both the amplitude square as well as the phase space considerations. Non-observation of any final particle is well taken care of by suitably performing corresponding phase space integration. We have already pointed out in the beginning of Sec. III of our paper [2] as well as in Sec.2.2 of [3] that such phase space integration for invisible neutrinos directly leads to pDMCT, and this does not require any ad hoc symmetrization of the Dirac case amplitude square as done in [1].
In the standard model (SM) the Dirac neutrino () and antineutrino () carry lepton numbers as they are produced via different weak charged currents (CCs) and for , to ensure lepton number conservation within the SM. When considering Dirac neutrino antineutrino pair production via two CCs, as it happens in the decays considered in Ref. [1] as well as in our work [2], the neutrino and antineutrino 4-momenta are paired with the corresponding charged lepton so that lepton number is conserved. This is not required for Majorana neutrinos, for which and the lepton number is not a good (conserved) quantum number. However, adding a term to the Dirac case amplitude square where the neutrino and antineutrino 4-momenta have been exchanged, as done in the right sides of Eqs. (16) and (35) of Ref. [1], amounts to considering and both of which violate lepton number. When working with Dirac neutrinos in the framework of the SM alone without consideration of any new physics contributions, as is the case in decay considered in both [1] and [2], one should avoid considering lepton number violation effects.
@@
We have considered the claims made in Ref. [1] in the context of our papers [2, 3, 4] and presented clarifications to consider them both invalid and incorrect. The main crux of the arguments presented in our papers [2, 3, 4] is to show that quantum statistics can not only distinguish between Majorana or Dirac types, but also define the domain of validity of pDMCT in the most general way possible in the absence of a formal proof of it until now. We note that pDMCT is strictly valid when we integrate over the momenta of neutrino and antineutrino (using the Dirac nomenclature). The proposed ad hoc symmetrisation in [1] at the probability level for the Dirac case simply achieves this by a different procedure, but as explained before, we do not consider this approach to be correct. Nevertheless, our main focus in [2, 3, 4, 5, 6] is on the possible exceptions to pDMCT when we can fix the neutrino-antineutrino momenta either directly (possible in principle but difficult experimentally) or indirectly (as pointed out in [2] under special kinematic conditions) or when new physics effects are taken into consideration.
@@ Acknowledgements.
The work of CSK is supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2022-NR070836, RS-2026-25471898). We thank the authors of [1], especially Innes Bigaran and Stephen Parke, for engaging in discussions on our paper that brings clarity to many of the issues.References
- [1] I. Bigaran, S. J. Parke and P. Pasquini, “Exploring Quantum Statistics for Dirac and Majorana Neutrinos using Spinor-Helicity techniques,” Phys. Rev. D 112, no.5, 053008 (2025) [arXiv:2507.07180 [hep-ph]].